How to solve: f(x)=(3x+2)/x
Inputs are: f(1), f(2)

Answer by stanbon(75887)

How to solve: f(x)=(3x+2)/x
Inputs are: f(1), f(2)
f(1) = (3*1+2)/1
f(1) = 5/1
f(1) = 5
--------------------
f(2) = (3*2+2)/2
f(2) = 8/2
f(2) = 4
=================
Cheers,
Stan H.